b) What is the Index of hydrogen deficiency of a compound with the following formula, CH.Br 3pts 2pts c) Give a possible stru please help
Answer #1

6 THO= 2(0)+2-(Pt) - = 2(5)+2-(7+) n = No. of carbons P = No of Hydrons = No. of bromine atoms - 1042 ام ۱۴۴ -(8) 16 18 IHO -